OpenStudy (anonymous):

did I do this problem right?Explain, in complete sentences, how you would completely factor 15x2 – 25x – 60 by using the grouping method. How would you check your factors for accuracy?

OpenStudy (anonymous):

First you would divide by 5: 5 (3x^3 - 5x - 12) now you need to find two numbers whose product is -12 and where the result of multiplying by 3 and adding give the middle number -5 - 3 * 4 = -12 and -3*3 + 4 = -5 so you then put them in parenthesis to get: (3x + 4)(x - 3) ( + 4x - 9x = -5x) then you get: 5(3x + 4)(x - 3) would this be right?

OpenStudy (anonymous):

no you are bot doing it right, you have to find two numbers whose product is =-12*3=-36 and not -12, so the numbers would be -9 and 4...now proceed using these valued to get a right answer

OpenStudy (anonymous):

3x²-9x+4x-12 (3x²-9x)+(4x-12) 3x(x-3)+4(x-3) x-3 is a common factor.so isolate that would be: (x-3)(3x+4) Solving: x = 3 or x = -4/3 {3, -4/3} would this be right ?

OpenStudy (anonymous):

yes u got it now

OpenStudy (anonymous):

ok thanks!!!